OpenStudy (geometry_hater):

Without graphing determine if the lines represented by the equation are same, are different and intersect at a point or are different and parallel x + 2y =8 2x + y = 4 this one is alot harder!

OpenStudy (jamesj):

Write both equations as y = something y = something else then use the same test we used last time

OpenStudy (geometry_hater):

wait what? im not following

OpenStudy (jamesj):

Last time you had lines in the form y = mx + c, where m and c are constants e.g., you had an equation like y = -4x + 3 We also last time had a test: If you have two lines in the form y = mx + c then they are parallel if and only if their slopes, the number m for both, are the same. We want now to use that same test here. But in order to do so, we need to write both lines in this form: y = mx + c. So take your two lines: x + 2y =8 2x + y = 4 and put them both in that form.

OpenStudy (geometry_hater):

ah it all makes sense now thank you!